### v1

@@ 1 Introduction
Correlation coefficients are fundamental tools for data analysis, quantifying the mutual dependence between two random variables. The classical and most popular correlation coefficients are Pearson correlation (pearson1896), Kendall’s $\tau$ (kendall1938) and Spearman’s $\rho$ (spearman1904), which indicate the direction of dependence by their sign, and the strength of dependence by their closeness to 1 in absolute value. Despite being the most widely used of the three coefficients, Pearson correlation, which we denote by $r$, has well-known shortcomings. It lacks fundamental theoretical properties such as invariance to monotonic transformations, making it dependent on data transformations, and attainability, in the sense that it can be far away from 1 in absolute value under very strong or perfect dependence, thereby compromising its interpretability (embrechts2002; fissler2023). Furthermore, it suffers from statistical problems in that its estimator is not robust to outliers and inefficient for heavy-tailed distributions. Finally, its applicability is limited: It is only defined for random variables on a metric scale, excluding ordinal variables, and only for variables with finite second moments, excluding heavy-tailed distributions. In the case of continuous random variables, the classical rank correlations Kendall’s $\tau$ and Spearman’s $\rho$ cure all of these shortcomings, but for discrete random variables, they also suffer from non-attainability (genest2007; neslehova2007). Popular generalizations of $\rho$ and $\tau$ mitigating this problem are Kendall’s $\tau_{b}$ (kendall1948) and grade correlation, which we denote by $\rho_{b}$. For $\tau$, there is even a simple and appealing generalization in Goodman-Kruskal’s $\gamma$ (goodmanKruskal1954), which fully cures this problem (pohle2025).
@@
This lack of key results is even more surprising given that hoeffding1948 in his seminal paper on U-statistics already in the iid case established asymptotic normality and derived the corresponding limiting variances of empirical Kendall’s $\tau$, $\widehat{\tau}$ (here even including the discrete case), and Spearman’s $\rho$, $\widehat{\rho}$ (limiting himself to the continuous case), making use of the facts that $\widehat{\tau}$ is a U-statistic and $\widehat{\rho}$ is asymptotically equivalent to a U-statistic. Also in the iid case, genest2013estimation provide the limiting distribution of multivariate generalizations of $\widehat{\rho}$, paying particular attention to the discrete case.
@@
In the time series case, dehling2017 derive an invariance principle for U-statistics of order 2 and in particular $\widehat{\tau}$ under weak dependence. The limiting distributions for the other rank correlations mentioned above, in particular $\widehat{\rho}$, do not even seem to be available in the continuous case from the extant literature.
@@
Moreover, estimators for asymptotic variances of rank correlations have been neglected in the literature with the consequence that classical tools for statistical inference – asymptotic confidence intervals and tests – are not available. The only exceptions are for Kendall’s $\tau$, where noether1967 discusses an estimator for the asymptotic variance (see also samara1988) and again dehling2017, who propose a variance estimator in the continuous time series case. Instead of consistent variance estimators, approximations to the asymptotic variances of $\widehat{\tau}$ and $\widehat{\rho}$ in the iid case seem to be in widespread use, see the proposal by fieller1957 being derived under bivariate normality, or modifications thereof (bonett2000). Already borkowf2002 demonstrated them to be inaccurate.
@@
The only settings, where classical asymptotic inference for $\tau$ and $\rho$ seems to be in widespread use, is in testing independence in the iid case. There, the limiting distributions under independence assuming continuous variables are used, where the limiting variances of $\widehat{\tau}$ and $\widehat{\rho}$ are $4/9$ and $1$, respectively. [^1] The basic R function cor.test from the stats package (RCoreTeam) only provides a confidence interval for Pearson correlation. For $\tau$ and $\rho$, only those tests of independence are available. The former result is due to kendall1938, and according to hoeffding1948, the latter is due to Student and was published by pearson1907. Thus, here, the variances are constant and do not have to be estimated. In the discrete case, however, the limiting variances under independence have not been available from the literature. hoeffding1948 conjectured that they are different from their values in the continuous case and depend on single-tie and double-tie probabilities of the two variables of interest.
@@
We provide a comprehensive treatment of asymptotic inference for classical rank correlations, namely for Kendall’s $\tau$ and Spearman’s $\rho$ as well as their generalizations to the discrete case, Goodman-Kruskal’s $\gamma$, $\tau_{b}$ and $\rho_{b}$. First, we establish asymptotic normality of their empirical versions, $\widehat{\tau}$, $\widehat{\rho}$, $\widehat{\gamma}$, $\widehat{\tau}_{b}$ and $\widehat{\rho}_{b}$, respectively, and derive the corresponding asymptotic variances under classical time series assumptions, which includes the important special case of iid processes. Here, we make use of asymptotic theory for U-statistics for weakly dependent stochastic processes (see dehling2006 for an overview). Secondly, we introduce consistent estimators for the asymptotic variances. Those two elements – the asymptotic distribution and consistent variance estimators – allow for the construction of confidence intervals and tests. Thirdly, we derive formulas for the asymptotic variances under independence between the two variables or processes of interest, which is important to execute widely-used tests of independence. In the iid case, our results generalize the aforementioned classical results from the continuous case. The variance formulas indeed depend on the double-tie probabilities and are easily implementable, providing simple and valid tests of independence also in the discrete case. In time series settings, our results generalize the results of lun2023, who establish the asymptotic distribution of $\widehat{\tau}$ and $\widehat{\rho}$ in the continuous case under independence between the two stochastic processes of interest. We also provide a multivariate central limit theorem for a vector of U-statistics, which enables the asymptotic analysis of $\widehat{\gamma}$, $\widehat{\tau}_{b}$ and $\widehat{\rho}_{b}$ as they are functions of several U-statistics. This result also demonstrates that vectors of empirical rank correlations are asymptotically multivariate normal and allows for the calculation of the corresponding covariance matrices.
@@
Appendix A contains a discussion of asymptotic theory for U-statistics under weak dependence and, in particular, the above-mentioned multivariate central limit theorem. Appendix B presents all proofs, and Appendix C an illustrative example involving a bivariate geometric distribution. An extensive description of the simulated DGPs from Section 6 and a detailed interpretation of the simulation results as well as the results itself in a tabulated form are shown in Appendix D. We also provide code implementing our confidence intervals and tests in an R package (RCoreTeam) at https://github.com/jan-lukas-wermuth/RCor and replication material at https://github.com/jan-lukas-wermuth/replication_RCor.
@@ 2.1 Fundamental Concepts: Grades, Sign Differences, and Tie Probabilities
We define the mid-distribution function (MDF) of $X$ as $G_{X}(x):=(F_{X}(x)+F_{X}(x^{-}))/2$ for $x\in\mathbb{R}$ (and analogously for $Y$), see e.g. parzen97 for a discussion. Note that the MDF can be rewritten as
@@
denotes the Heaviside step function using the half-maximum convention. Essentially, the MDF counts the probability mass for values smaller than $x$ fully, and the probability mass for values equal to $x$ half. It is thus closely related to the concept of a midrank, see (11) in the next section. Furthermore, we define the bivariate MDF of $(X,Y)$ as (see e.g. hoeffding1948)
@@
Plugging-in a univariate or bivariate random variable itself into the respective MDF, we obtain the univariate grade (hoeffding1948) or mid-distribution transform (parzen2004), $G_{X}(X)$, and the bivariate grade $G_{X,Y}(X,Y)$, which are of fundamental importance for this paper. The univariate grade is a variant of the probability integral transform (PIT) $F_{X}(X)$ and better-suited for arbitrary real-valued random variables. It reduces to the PIT if $F_{X}$ is continuous. While the univariate grade does not follow a standard uniform distribution in the discrete case like the PIT (and itself) in the continuous case ($F_{X}(X)\sim U(0,1)$ if $X$ is continuous), contrary to the PIT its expectation is always $1/2$. Also, the variance is smaller than $1/12$ by a factor only depending on the double-tie probability of $X$, i.e. on $\mathrm{P}(X=X^{\prime}=X^{\prime\prime})$, also see the subsequent Lemma 2.1. For the moments of the bivariate grade, there are no general results available, and its distribution depends on the distribution of $(X,Y)$. We often need its expectation in the paper, which we consequently need to estimate.
@@ 2.2 Classical Rank Correlations: Kendall’s $\tau$ and Spearman’s $\rho$
By part (iii) of Lemma 2.1, it holds that $\tau(X,Y)=\mathrm{Cov}\big(\mathrm{sgn}(X-X^{\prime}),\mathrm{sgn}(Y-Y^{\prime})\big)$, and in the continuous case, we even have $\tau(X,Y)=\mathrm{Cor}\big(\mathrm{sgn}(X-X^{\prime}),\mathrm{sgn}(Y-Y^{\prime})\big)$. Kendall’s $\tau$ can also be compactly written in terms of the grade by using parts (ii) and (iv) of Lemma 2.1, namely $\tau(X,Y)=4\mathrm{E}[G_{X,Y}(X,Y)]-1$, which generalizes its relation to the joint CDF and copula from the continuous case (schweizer1981). For the special case of the bivariate geometric distribution, see Appendix C.
@@
A suitable population version of Spearman’s $\rho$ involves the covariance of the grades, $G_{X}(X)$ and $G_{Y}(Y)$, see e.g. hoeffding1948.
@@
Kendall’s $\tau$ and Spearman’s $\rho$ possess all major desirable properties for dependence measures in the continuous case (embrechts2002). However, for discrete random variables, they lack the property of attainability (neslehova2007; genest2007), that is, they do not take the values 1-1 and $1$ under perfect negative and positive dependence (counter- and comonotonicity), respectively. In fact, they may be very far away from those values under perfect or strong dependence. This leads to problems with their interpretability as strong dependence should be indicated by values close to 1 in absolute value. pohle2025 analyze attainable generalizations of $\tau$ and $\rho$, which thus fulfill all desirable properties in general, and refer to them as “proper”. For Kendall’s $\tau$, there exists a particularly nice and simple proper counterpart, namely Goodman-Kruskal’s $\gamma$. This measure just computes Kendall’s $\tau$ conditional on the event that there are no ties in both $X$ and $Y$, recall (5).
@@
However, while $\tau_{b}$ mitigates the problem in that it is weakly larger in absolute value than $\tau$, pohle2025 show that it is still non-attainable and thus inferior to $\gamma$. Nevertheless, we consider $\tau_{b}$ due to its popularity and the widespread use of its empirical analogue in statistical software.
@@ 3.1 Classical Empirical Rank Correlations
where $k^{(\rho)}$ is defined in (2.2). More precisely, the relation between $\widehat{\rho}$ and $\widetilde{\rho}$ is the following, see hoeffding1948, which arises essentially by the relation between midranks and the difference sign function from (11): [^2] The slightly different factors involving $n$ in hoeffding1948 arise because Hoeffding defines empirical Spearman’s $\rho$ as $n^{2}/(n^{2}-1)\,\widehat{\rho}$.
@@ 4.1 Consistency
Note that we do not need moment conditions as the considered rank-based measures do not operate with the observations themselves, but with their ranks, grades, or CDFs, respectively. All limit theorems from this section are special cases of the more general results in the next section, where we work under classical time series assumptions, so that the formulas for the asymptotic variances get more involved. We nevertheless dedicate a separate section to the results for iid data because they are of high relevance in their own right and new to a large part. Furthermore, it is beneficial to appreciate the less involved results in this section, in particular regarding variance estimation, before tackling the cumbersome time series results. In Appendix A, we review a univariate central limit theorem (Proposition A.2) for U-statistics under the iid assumption (hoeffding1948) and under weak dependence (denker1983). Furthermore, we present a multivariate central limit theorem for U-statistics (Proposition A.3) under both assumptions, which in combination with the delta method (see beutner24) allows for the treatment of $\widehat{\gamma}$, $\widehat{\tau}_{b}$, and $\widehat{\rho}_{b}$.
The following consistency result follows from the classical law of large numbers for U-statistics (hoeffding1948) and the continuous mapping theorem.
@@ 4.2 Asymptotic Distributions
and the latter is computed by part (ii) of Lemma 2.1. The full derivations of the other leading terms appearing in Proposition 4.2 and 4.3 are provided by Lemma B.2 in the appendix. The asymptotic distribution of $\widehat{\tau}$ has already been established by hoeffding1948 and, in the continuous case, also the asymptotic distribution of $\widehat{\rho}$.

### v2

@@ 1 Introduction
Correlation coefficients are fundamental tools for data analysis, quantifying the mutual dependence between two random variables. The classical and most popular correlation coefficients are Pearson correlation (Pearson, 1896), Kendall’s $\tau$ (Kendall, 1938) and Spearman’s $\rho$ (Spearman, 1904), which indicate the direction of dependence by their sign, and the strength of dependence by their closeness to 1 in absolute value. Despite being the most widely used of the three coefficients, Pearson correlation, which we denote by $r$, has well-known shortcomings. It lacks fundamental theoretical properties such as invariance to monotonic transformations, making it dependent on data transformations, and attainability, in the sense that it can be far away from 1 in absolute value under very strong or perfect dependence, thereby compromising its interpretability (Embrechts et al., 2002; Fissler and Pohle, 2023). Furthermore, it suffers from statistical problems in that its estimator is not robust to outliers and inefficient for heavy-tailed distributions. Finally, its applicability is limited: It is only defined for random variables on a metric scale, excluding ordinal variables, and only for variables with finite second moments, excluding heavy-tailed distributions. In the case of continuous random variables, the classical rank correlations Kendall’s $\tau$ and Spearman’s $\rho$ cure all of these shortcomings, but for discrete random variables, they also suffer from non-attainability (Genest and Nešlehová, 2007; Nešlehová, 2007). Popular generalizations of $\rho$ and $\tau$ mitigating this problem are Kendall’s $\tau_{b}$ (Kendall, 1948) and grade correlation, which we denote by $\rho_{b}$. For $\tau$, there is even a simple and appealing generalization in Goodman-Kruskal’s $\gamma$ (Goodman and Kruskal, 1954), which fully cures this problem (Pohle and Wermuth, 2025).
@@
This lack of key results is even more surprising given that Hoeffding (1948) in his seminal paper on U-statistics already in the iid case established asymptotic normality and derived the corresponding limiting variances of empirical Kendall’s $\tau$, $\widehat{\tau}$ (here even including the discrete case), and Spearman’s $\rho$, $\widehat{\rho}$ (limiting himself to the continuous case), making use of the facts that $\widehat{\tau}$ is a U-statistic and $\widehat{\rho}$ is asymptotically equivalent to a U-statistic. Also in the iid case, Genest et al. (2013) provide the limiting distribution of multivariate generalizations of $\widehat{\rho}$, paying particular attention to the discrete case.
@@
In the time series case, Dehling et al. (2017) derive an invariance principle for U-statistics of order 2 and in particular $\widehat{\tau}$ under weak dependence. The limiting distributions for the other rank correlations mentioned above, in particular $\widehat{\rho}$, do not even seem to be available in the continuous case from the extant literature.
@@
Moreover, estimators for asymptotic variances of rank correlations have been neglected in the literature with the consequence that classical tools for statistical inference – asymptotic confidence intervals and tests – are not available. The only exceptions are for Kendall’s $\tau$, where Noether (1967) discusses an estimator for the asymptotic variance (see also Samara and Randles, 1988) and again Dehling et al. (2017), who propose a variance estimator in the continuous time series case. Instead of consistent variance estimators, approximations to the asymptotic variances of $\widehat{\tau}$ and $\widehat{\rho}$ in the iid case seem to be in widespread use, see the proposal by Fieller et al. (1957) being derived under bivariate normality, or modifications thereof (Bonett and Wright, 2000). Already Borkowf (2002) demonstrated them to be inaccurate.
@@
The only settings, where classical asymptotic inference for $\tau$ and $\rho$ seems to be in widespread use, is in testing independence in the iid case. There, the limiting distributions under independence assuming continuous variables are used, where the limiting variances of $\widehat{\tau}$ and $\widehat{\rho}$ are $4/9$ and $1$, respectively. [^1] The basic R function cor.test from the stats package (R Core Team, 2025) only provides a confidence interval for Pearson correlation. For $\tau$ and $\rho$, only those tests of independence are available. The former result is due to Kendall (1938), and according to Hoeffding (1948, p. 321), the latter is due to Student and was published by Pearson (1907). Thus, here, the variances are constant and do not have to be estimated. In the discrete case, however, the limiting variances under independence have not been available from the literature. Hoeffding (1948) conjectured that they are different from their values in the continuous case and depend on single-tie and double-tie probabilities of the two variables of interest.
@@
We provide a comprehensive treatment of asymptotic inference for classical rank correlations, namely for Kendall’s $\tau$ and Spearman’s $\rho$ as well as their generalizations to the discrete case, Goodman-Kruskal’s $\gamma$, $\tau_{b}$ and $\rho_{b}$. First, we establish asymptotic normality of their empirical versions, $\widehat{\tau}$, $\widehat{\rho}$, $\widehat{\gamma}$, $\widehat{\tau}_{b}$ and $\widehat{\rho}_{b}$, respectively, and derive the corresponding asymptotic variances under classical time series assumptions, which includes the important special case of iid processes. Here, we make use of asymptotic theory for U-statistics for weakly dependent stochastic processes (see Dehling (2006) for an overview). Secondly, we introduce consistent estimators for the asymptotic variances. Those two elements – the asymptotic distribution and consistent variance estimators – allow for the construction of confidence intervals and tests. Thirdly, we derive formulas for the asymptotic variances under independence between the two variables or processes of interest, which is important to execute widely-used tests of independence. In the iid case, our results generalize the aforementioned classical results from the continuous case. The variance formulas indeed depend on the double-tie probabilities and are easily implementable, providing simple and valid tests of independence also in the discrete case. In time series settings, our results generalize the results of Lun et al. (2023), who establish the asymptotic distribution of $\widehat{\tau}$ and $\widehat{\rho}$ in the continuous case under independence between the two stochastic processes of interest. We also provide a multivariate central limit theorem for a vector of U-statistics, which enables the asymptotic analysis of $\widehat{\gamma}$, $\widehat{\tau}_{b}$ and $\widehat{\rho}_{b}$ as they are functions of several U-statistics. This result also demonstrates that vectors of empirical rank correlations are asymptotically multivariate normal and allows for the calculation of the corresponding covariance matrices.
@@
Appendix A contains a discussion of asymptotic theory for U-statistics under weak dependence and, in particular, the above-mentioned multivariate central limit theorem. Appendix B presents all proofs, and Appendix C an illustrative example involving a bivariate geometric distribution. An extensive description of the simulated DGPs from Section 6 and a detailed interpretation of the simulation results as well as the results itself in a tabulated form are shown in Appendix D. We also provide code implementing our confidence intervals and tests in an R package (R Core Team, 2025) at https://github.com/jan-lukas-wermuth/RCor and replication material at https://github.com/jan-lukas-wermuth/replication_RCor.
@@ 2.1 Fundamental Concepts: Grades, Sign Differences, and Tie Probabilities
We define the mid-distribution function (MDF) of $X$ as $G_{X}(x):=(F_{X}(x)+F_{X}(x^{-}))/2$ for $x\in\mathbb{R}$ (and analogously for $Y$), see e.g. Parzen (1997) for a discussion. Note that the MDF can be rewritten as
@@
denotes the Heaviside step function using the half-maximum convention. Essentially, the MDF counts the probability mass for values smaller than $x$ fully, and the probability mass for values equal to $x$ half. It is thus closely related to the concept of a midrank, see (11) in the next section. Furthermore, we define the bivariate MDF of $(X,Y)$ as (see e.g. Hoeffding, 1948)
@@
Plugging-in a univariate or bivariate random variable itself into the respective MDF, we obtain the univariate grade (Hoeffding, 1948) or mid-distribution transform (Parzen, 2004), $G_{X}(X)$, and the bivariate grade $G_{X,Y}(X,Y)$, which are of fundamental importance for this paper. The univariate grade is a variant of the probability integral transform (PIT) $F_{X}(X)$ and better-suited for arbitrary real-valued random variables. It reduces to the PIT if $F_{X}$ is continuous. While the univariate grade does not follow a standard uniform distribution in the discrete case like the PIT (and itself) in the continuous case ($F_{X}(X)\sim U(0,1)$ if $X$ is continuous), contrary to the PIT its expectation is always $1/2$. Also, the variance is smaller than $1/12$ by a factor only depending on the double-tie probability of $X$, i.e. on $\mathrm{P}(X=X^{\prime}=X^{\prime\prime})$, also see the subsequent Lemma 2.1. For the moments of the bivariate grade, there are no general results available, and its distribution depends on the distribution of $(X,Y)$. We often need its expectation in the paper, which we consequently need to estimate.
@@ 2.2 Classical Rank Correlations: Kendall’s $\tau$ and Spearman’s $\rho$
By part (iii) of Lemma 2.1, it holds that $\tau(X,Y)=\mathrm{Cov}\big(\mathrm{sgn}(X-X^{\prime}),\mathrm{sgn}(Y-Y^{\prime})\big)$, and in the continuous case, we even have $\tau(X,Y)=\mathrm{Cor}\big(\mathrm{sgn}(X-X^{\prime}),\mathrm{sgn}(Y-Y^{\prime})\big)$. Kendall’s $\tau$ can also be compactly written in terms of the grade by using parts (ii) and (iv) of Lemma 2.1, namely $\tau(X,Y)=4\mathrm{E}[G_{X,Y}(X,Y)]-1$, which generalizes its relation to the joint CDF and copula from the continuous case (Schweizer and Wolff, 1981). For the special case of the bivariate geometric distribution, see Appendix C.
@@
A suitable population version of Spearman’s $\rho$ involves the covariance of the grades, $G_{X}(X)$ and $G_{Y}(Y)$, see e.g. Hoeffding (1948).
@@
Kendall’s $\tau$ and Spearman’s $\rho$ possess all major desirable properties for dependence measures in the continuous case (Embrechts et al., 2002). However, for discrete random variables, they lack the property of attainability (Nešlehová, 2007; Genest and Nešlehová, 2007), that is, they do not take the values 1-1 and $1$ under perfect negative and positive dependence (counter- and comonotonicity), respectively. In fact, they may be very far away from those values under perfect or strong dependence. This leads to problems with their interpretability as strong dependence should be indicated by values close to 1 in absolute value. Pohle and Wermuth (2025) analyze attainable generalizations of $\tau$ and $\rho$, which thus fulfill all desirable properties in general, and refer to them as “proper”. For Kendall’s $\tau$, there exists a particularly nice and simple proper counterpart, namely Goodman-Kruskal’s $\gamma$. This measure just computes Kendall’s $\tau$ conditional on the event that there are no ties in both $X$ and $Y$, recall (5).
@@
However, while $\tau_{b}$ mitigates the problem in that it is weakly larger in absolute value than $\tau$, Pohle and Wermuth (2025) show that it is still non-attainable and thus inferior to $\gamma$. Nevertheless, we consider $\tau_{b}$ due to its popularity and the widespread use of its empirical analogue in statistical software.
@@ 3.1 Classical Empirical Rank Correlations
where $k^{(\rho)}$ is defined in (2.2). More precisely, the relation between $\widehat{\rho}$ and $\widetilde{\rho}$ is the following, see Hoeffding (1948, eq. (9.21)), which arises essentially by the relation between midranks and the difference sign function from (11): [^2] The slightly different factors involving $n$ in Hoeffding (1948, eq. (9.21)) arise because Hoeffding defines empirical Spearman’s $\rho$ as $n^{2}/(n^{2}-1)\,\widehat{\rho}$.
@@ 4.1 Consistency
Note that we do not need moment conditions as the considered rank-based measures do not operate with the observations themselves, but with their ranks, grades, or CDFs, respectively. All limit theorems from this section are special cases of the more general results in the next section, where we work under classical time series assumptions, so that the formulas for the asymptotic variances get more involved. We nevertheless dedicate a separate section to the results for iid data because they are of high relevance in their own right and new to a large part. Furthermore, it is beneficial to appreciate the less involved results in this section, in particular regarding variance estimation, before tackling the cumbersome time series results. In Appendix A, we review a univariate central limit theorem (Proposition A.2) for U-statistics under the iid assumption (Hoeffding, 1948) and under weak dependence (Denker and Keller, 1983). Furthermore, we present a multivariate central limit theorem for U-statistics (Proposition A.3) under both assumptions, which in combination with the delta method (see Beutner, 2024) allows for the treatment of $\widehat{\gamma}$, $\widehat{\tau}_{b}$, and $\widehat{\rho}_{b}$.
The following consistency result follows from the classical law of large numbers for U-statistics (Hoeffding, 1948) and the continuous mapping theorem.
@@ 4.2 Asymptotic Distributions
and the latter is computed by part (ii) of Lemma 2.1. The full derivations of the other leading terms appearing in Proposition 4.2 and 4.3 are provided by Lemma B.2 in the appendix. The asymptotic distribution of $\widehat{\tau}$ has already been established by Hoeffding (1948) and, in the continuous case, also the asymptotic distribution of $\widehat{\rho}$.